### v1

@@ 3.2 Effect of Intrinsic Plane Height
Figure 4: Effect of proper motion uncertainties on satellite galaxies under different initial vertical distributions of $h=$ 10, $h$ = 20 kpc and $h$= 30 kpc. The left panel shows proper motion uncertainties of 0.04 mas yr-1, while the right panel displays uncertainties of 0.08 mas yr-1. Dashed curves indicate forward integration, and solid curves represent backward integration.
We also tested whether the effect of proper motion uncertainties on the plane of satellite galaxies is sensitive to our choice of the initial vertical distribution of the intrinsic satellite planes. Specifically, we ran the system with three different plane heights — $h=10\,\text{kpc}$, $h=20\,\text{kpc}$, and $h=30\,\text{kpc}$ — bracketing the range of observationally inferred satellite plane heights for the Milky Way and M31 (Ibata et al., 2013; Pawlowski et al., 2013).
Fig. 4 displays two panels showing the results for adopted proper motion uncertainties of 0.04 mas yr-1 and 0.08 mas yr-1. In both plots, the dotted-dashed curves represent mean axis ratios of the forward integrations for the different heights of $h=10,20,$ and $30$ kpc, indicated by orange, gray, and cyan colors, respectively. As expected, when the vertical distribution is lower, the mean axis ratio of the system is lower, making the system flatter compared to the other values of $h$. In the backward integration, after adding uncertainties in proper motion, we observe that all three curves exhibit a similar trend: errors in proper motion lead to an apparent widening of the satellite plane. This effect is substantial in all three cases, indicating that our findings are not sensitive to the exact choice of intrinsic plane height.
Table 3 summarizes the quantitative impact of proper motion uncertainties on the inferred change in the mean axis ratio. For a smaller intrinsic plane height of $h=10\,$ kpc (model BI-L10 and BI-H10 in Table 3), the relative effect is stronger, with $f_{c/a}=$ 2.45 (3.94) compared to $f_{c/a}=$ 1.64 (2.25) for the model of $h=$ 20 kpc for adopted proper motion uncertainties of 0.04 (0.08) mas yr-1. Similarly, for an intrinsically wider plane (model BI-L30 and BI-H30 in Table 3), the relative change in plane flattening is a bit lower, with $f_{c/a}=$ 1.29 and 1.72 for proper motion uncertainties of 0.04 and 0.08 mas yr-1, respectively. However, the absolute change in axis ratio, measured via $\Delta_{c/a}$, does not differ as much.
Table 3: Parameters and results for backward integration models at 3 Gyr for different intrinsic plane height, similar to Table 4.
@@ 3.3 Effect of Distance Uncertainties
Figure 5: Similar to Fig. 3, but here we compare it with 5% distance uncertainties in the system as well. The green curve in each panel shows 0% distance uncertainties, whereas the red curve shows 5% distance uncertainties. The uncertainties in proper motions are as follows: 0.00 mas yr-1 in the upper left panel, 0.04 mas yr-1 in the upper right panel, 0.08 mas yr-1 in the lower left panel, and 0.12 mas yr-1 in the lower right panel. Both forward and backward integrations are done using the $MW_{\text{fiducial}}$ potential model.
We now go one step further and additionally introduce 5% distance uncertainties before backwards-integrating, to study the behavior of the plane under both proper motions and distance uncertainties. In all four panels of Fig. 5, backward integration now has two curves: one without distance uncertainties (shown in green) and one with 5% distance uncertainties (shown in red). As shown in the upper left panel of Fig. 5, when proper motion uncertainties are ignored, there is a noticeable effect of distance uncertainties; the system does not return to its original state but instead experiences a slight change. Despite this, there is still decent agreement between the mean backward and forward integrations in axis ratio, suggesting that the inferred fundamental orbital evolution is not drastically impacted, even with the introduction of 5% uncertainties in the distances of the test satellites. It is also evident from Table 4 where in models BI-01 and B1-13, both $\Delta_{c/a}$ and $f_{c/a}$ have changed just slightly from 0.00 to 0.02 and 1.00 to 1.17 respectively after 3 Gyr of backward integration. A 5% error in the distances thus translates to a widening of the inferred past plane height over the true one, reaching up to 17% at specific times (e.g., after 3 Gyr), although the deviation remains negligible for the majority of the 5 Gyr integration period.
@@
We also conducted additional tests where we examined the correlation between proper motion errors and distance errors. We found that although there is a noticeable change, the overall effect is not extreme. This is discussed in Appendix B.1.
@@ 3.4 Effect of Adopted Potential
Figure 6: Evolution of the plane of satellite galaxies with proper motion uncertainties included in each panel. The black dashed-dotted curves represent the mean axis ratio for forward integration, while the colored dashed curves indicate the mean axis ratio for backward integration under different potential models characterized by halo mass. Specifically, the blue, green, and red dashed curves correspond to the ${c/a}$ values under the $MW_{\text{low}}$, $MW_{\text{fiducial}}$, and $MW_{\text{high}}$ potential models, respectively. Note that these results do not account for distance uncertainties.
@@
In the upper left panel of Fig. 6, without the uncertainties in proper motion and distance, the results provide a clear picture of the satellite plane’s fundamental dynamical behavior when subjected to different potential models. The colored curves depict the mean backward integration under different Milky Way potentials. The backward integration, particularly the green curve representing the $MW_{\text{fiducial}}$ potential, demonstrates perfect consistency with its forward counterpart, implying a stable evolution when the potential model is the same. This case is already discussed in the previous sections.
@@
The red curve, representing the $MW_{\text{high}}$ model, exhibits a reasonable degree of stability, although it does not completely resemble the forward integration, suggesting some sensitivity to the host galaxy’s mass assumptions. This difference is shown in BI-09 of Table 4, which shows $\Delta_{c/a}$ = 0.01 after 3 Gyr, and $f_{c/a}$ = 1.13. In contrast, the blue curve, representing the $MW_{\text{low}}$ mass model, shows a completely different behavior from the other models. Its axis ratio initially increases before flattening, which could imply a dynamical history distinct from the correct one for the plane of satellites within such a lower-mass potential model. Here, for lower-mass Milky Way potential, $\Delta_{c/a}$ = 0.03 after 3 Gyr, and $f_{c/a}$ = 1.27.
@@
The other three panels of Fig. 6 show the effect of both potential models and proper motion uncertainties on the stability of the plane of satellites. In the upper right panel, we have proper motion uncertainties of $\pm$0.04 mas yr-1. The effect under $MW_{\text{fiducial}}$ has already been discussed in the previous section. For the $MW_{\text{high}}$ model, it generally follows a similar trend to $MW_{\text{fiducial}}$, but with slightly larger axis ratio values. Whereas, the $MW_{\text{low}}$ again shows a distinct behavior where it initially rises and then flattens out over time.
@@
When adding more uncertainties, up to Gaia levels and even with some extremes, the effect strengthens further in all the remaining panels, and the plane of the system becomes even wider. Under 0.12 mas yr-1 proper motion uncertainty, all three systems seem to have a similar trend, suggesting that higher proper motion uncertainties overshadow the effect of using an incorrect Milky Way model. This observation indicates that, at higher levels of proper motion uncertainty, the distinctions between the various Milky Way potential models become less pronounced, as the increased uncertainty tends to dominate the inferred dynamics of the satellite plane.
@@
Building upon this, we add one more layer: distance uncertainties. By comparing Fig. 6 and Fig. 7, we can again see a noticeable effect of distance uncertainties on the system under all three potential models, but it does not change the evolution of the plane drastically. However, in the presence of higher proper motion uncertainties, the effect of distance uncertainties is again overshadowed under all three potential models. Models BI-13 to BI-24 show this behavior when 5% of uncertainties were induced.
@@
We also checked that our results remain qualitatively unchanged when using different Milky Way gravitational potentials. Specifically, we performed the analysis with MilkyWayPotential2022 from Gala (Price-Whelan, 2017), including variations in halo mass and proper motion uncertainties, and found no substantial differences in the overall trends; see Appendix B.2.
@@ 3.5 Analysis of Individual Axes
Figure 8: Comparative analysis of galactic orbital axes under proper motion uncertainties and different potential models at $\theta_{tan}$ = 80°. It shows the mean values of 600 major, minor, and intermediate axes calculated through backward integration. Each panel represents a unique combination of proper motion uncertainty and potential models. Dotted-dashed curves show the effect of distance uncertainties, while solid curves ignores it. The rows correspond to different levels of proper motion uncertainty, while the columns represent the three axes: major (first column), intermediate (second column), and minor (third column).
@@
The behavior of the minor-to-major axial ratio in the $MW_{\text{low}}$ potential model differs from other models, warranting a closer examination of how the extent of the test satellite system changes along its different axes, from major to minor. Fig. 8 illustrates the influence of proper motion uncertainties on the mean values of the three axes: $a$ (major axis), $b$ (intermediate axis), and $c$ (minor axis). The influence of proper motion uncertainties is presented row-wise, while the axes are displayed column-wise in the figure. Each panel further shows curves in three colors, depicting the impact of potential models. The solid curves represent the results when $\epsilon_{dist}=0\%$, while the dashed-dotted curves indicate the results when $\epsilon_{dist}=5\%$. The first row shows $\pm$0.00 mas yr-1 proper motion uncertainties and 0% distance uncertainties. Here, both $MW_{\text{fiducial}}$ and $MW_{\text{high}}$ show similar behavior under respective axes, with and without distance uncertainties. Over time, the axes tend to decrease, particularly under $MW_{\text{high}}$ potential model. In contrast, $MW_{\text{low}}$ demonstrates a completely different and opposite behavior. Over time, in backward integration, all the axes under $MW_{\text{low}}$ tend to increase.
@@
In these models, the variations in axis lengths are a direct consequence of changes in gravitational pull due to mass differences. In the $MW_{\text{high}}$ model, with increased halo mass, all axes become more compact as the structure becomes more gravitationally bound. Conversely, in the $MW_{\text{low}}$ model, a reduction in mass leads to an elongation of the axes, particularly in $a$ and $b$. This suggests that in the $MW_{\text{low}}$ model, more satellites become unbound, causing them to drift away from their host galaxies. It is important to note that the $MW_{\text{fiducial}}$ model was used as a baseline for the $MW_{\text{high}}$ and $MW_{\text{low}}$ models, with the halo mass being adjusted by 40% (increased and decreased, respectively). However, the response of the axis extents to these adjustments is striking, with the Milky Way potential with 40% less halo mass having a more pronounced effect on the system than the one with a 40% higher halo mass. Moreover, with the inclusion of distance errors, each axis in every potential model extends slightly, indicating a noticeable, though not extreme, effect on the orbital axes.
@@
We additionally tested and recalculated the $c/a$ ratio by excluding test satellites beyond 300 kpc and only considering those within this threshold, as discussed in Appendix B.3. We find that this does not result in any major influence.
@@ 3.6 Inferred Escaping Satellites
Figure 9: Average number of test satellite exceeding 300 kpc of radial distance under proper motion uncertainties. The four panels represent different level of proper motion uncertainties. Each panel shows six curves, which visualize the number of test satellites residing beyond 300 kpc of distance under three different potential models. Here dotted-dashed curves show the results when including distance uncertainties, while solid line ignore distance uncertainties and only consider the proper motion errors.
Fig. 9 shows the average number of test satellites that surpass a distance of 300 kpc from the galactic center under proper motion uncertainties and various potential models. This plot is divided into four panels, and each panel further shows curves in three colors, depicting the impact of potential models. The solid curves represent the results when $\epsilon_{dist}=0\%$, while the dashed-dotted curves indicate the results when $\epsilon_{dist}=5\%$. In the upper left panel, when there are no proper motion uncertainties, for $MW_{\text{low}}$, the curve initially shows a relatively stable number of satellites that go beyond 300 kpc in backward integration, with and without distance uncertainties. Over time, the number gradually increases before reaching a plateau. This outcome is expected, as reducing the halo mass by 40% lowers the gravitational binding force, allowing more satellites to become unbound and drift beyond 300 kpc.
Unlike $MW_{\text{low}}$ model, the other two potential models show an opposite behavior. In $MW_{\text{fiducial}}$, the number of test satellites going beyond 300 kpc decreases over time, and at 5 Gyr, all $N_{\text{sat}}$ test satellites remain under the 300 kpc limit. This trend is even more pronounced in the $MW_{\text{high}}$ potential model, where the number of distant test satellites progressively diminishes over time due to the higher halo mass. Furthermore, the other three panels demonstrate the effect of proper motion uncertainties on satellite dynamics, showing that as uncertainties increase, more test satellites remain at distances greater than 300 kpc across all models. Proper motion errors affect this on average by preferentially increasing the inferred velocities and thus the energies of satellites, effectively ejecting more test satellites.
Furthermore, when comparing the curves with and without distance uncertainties, we observe a clear effect of these uncertainties, which further increases the number of test satellites that exceed the 300 kpc range. In Table 4, the column $N_{\text{sat}}>$ 300 kpc represents models BI-05 to BI-08 without induced distance errors in $MW_{\text{low}}$, and models BI-17 to BI-20 with distance errors. From this, it is evident that the inclusion of distance errors at 3 Gyr leads to a greater inferred loss of satellites in the system.

### v2

@@ 3.2 Effect of intrinsic plane height
Figure 4: Effect of proper motion uncertainties on satellite galaxies under different initial vertical distributions of $h=$ 10, $h$ = 20 kpc, and $h$= 30 kpc. The left panel shows proper motion uncertainties of 0.04 mas yr-1, while the right panel displays uncertainties of 0.08 mas yr-1. Dashed curves indicate forward integrations and solid curves represent backward integrations.
We also tested whether the effect of proper motion uncertainties on the plane of satellite galaxies was sensitive to our choice of the initial vertical distribution of the intrinsic satellite planes. Specifically, we ran the system with three different plane heights — $h=10\,\text{kpc}$, $h=20\,\text{kpc}$, and $h=30\,\text{kpc}$ — bracketing the range of observationally inferred satellite plane heights for the Milky Way and M31 (Ibata et al., 2013; Pawlowski et al., 2013).
Figure 4 displays two panels showing the results for adopted proper motion uncertainties of 0.04 mas yr-1 and 0.08 mas yr-1. In both plots, the dash-dotted curves represent mean axis ratios of the forward integrations for different heights of $h=10,20,$ and $30$ kpc, indicated by orange, gray and cyan, respectively. As expected, when the vertical distribution is lower, the mean axis ratio of the system is lower, making the system flatter compared to the other values of $h$. In backward integration, after adding proper motion uncertainties, we observe that all three curves exhibit a similar trend: errors in proper motion lead to an apparent widening of the satellite plane. This effect is substantial in all three cases, indicating that our findings are insensitive to the exact choice of intrinsic plane height.
Table 3 summarizes the quantitative impact of proper motion uncertainties on the inferred change in the mean axis ratio. For a lower intrinsic plane height of $h=10\,$ kpc (model BI-L10 and BI-H10 in Table 3), the relative effect is stronger, with $f_{c/a}=$ 2.45 (3.94) compared to $f_{c/a}=$ 1.64 (2.25) for the model with $h=$ 20 kpc, for adopted proper motion uncertainties of 0.04 (0.08) mas yr-1. Similarly, for an intrinsically wider plane (models BI-L30 and BI-H30 in Table 3), the relative change in plane flattening is slightly lower, with $f_{c/a}=$ 1.29 and 1.72 for proper motion uncertainties of 0.04 and 0.08 mas yr-1, respectively. However, the absolute change in axis ratio, measured via $\Delta_{c/a}$, does not differ as significantly.
Table 3: Parameters and results for backward integration models at 3 Gyr for different intrinsic plane heights, similar to Table 4
@@ 3.3 Effect of distance uncertainties
Figure 5: Similar to Fig. 3, but including a comparison with 5% distance uncertainties in the system. The green curve in each panel represents 0% distance uncertainties, while the red curve shows 5%. Proper motion uncertainties are: 0.00 mas yr-1 (upper left), 0.04 mas yr-1 (upper right), 0.08 mas yr-1 (lower left), and 0.12 mas yr-1 (lower right). Both forward and backward integrations are performed using the $MW_{\text{fiducial}}$ potential model.
To further assess the impact of observational uncertainties, we introduced additional 5% distance uncertainties prior to backward integration. This allowed us to study the behavior of the plane under both proper motion and distance uncertainties. In all four panels of Fig. 5, backward integration includes two curves: one without distance uncertainties (shown in green) and one with 5% distance uncertainties (shown in red). As shown in the upper left panel of Fig. 5, when proper motion uncertainties are ignored, distance uncertainties have a noticeable effect: the system does not return to its original state but instead experiences a slight change. Despite this, the mean axis ratios from the backward and forward integrations remain in good agreement, suggesting that the inferred fundamental orbital evolution is not significantly affected, even with the introduction of 5% uncertainties in the distances of the test satellites. This is also evident in Table 4, where in models BI-01 and BI-13, both $\Delta_{c/a}$ and $f_{c/a}$ change only slightly after 3 Gyr of backward integration, from 0.00 to 0.02 and from 1.00 to 1.17, respectively. Thus, a 5% error in the distances translates to a widening of the inferred past plane height over the true value, reaching up to 17% at specific times (e.g., after 3 Gyr), although the deviation remains negligible for the majority of the 5 Gyr integration period.
@@
We also performed additional tests to examine the correlation between proper motion errors and distance errors. We find that, although there is a noticeable change, the overall effect is not extreme. This is discussed in Appendix B.1.
@@ 3.4 Effect of adopted potential
Figure 6: Evolution of the plane of satellite galaxies with proper motion uncertainties included in each panel. The black dashed-dotted curves represent the mean axis ratio for forward integration, while the colored dashed curves show the mean axis ratio for backward integration under different potential models characterized by halo mass. Specifically, the blue, green, and red dashed curves correspond to the ${c/a}$ values under the $MW_{\text{low}}$, $MW_{\text{fiducial}}$, and $MW_{\text{high}}$ potential models, respectively. Note that these results do not account for distance uncertainties.
@@
In the upper left panel of Fig. 6, without the uncertainties in proper motion and distance, the results provide a clear picture of the fundamental dynamical behavior of the satellite plane when subjected to different potential models. The colored curves depict the mean backward integration under different Milky Way potentials. The backward integration, particularly the green curve representing the $MW_{\text{fiducial}}$ potential, demonstrates perfect consistency with its forward counterpart, implying a stable evolution when the potential model is the same. This case has been discussed in the previous sections.
@@
The red curve, representing the $MW_{\text{high}}$ model, exhibits a reasonable degree of stability, although it does not completely resemble the forward integration, suggesting some sensitivity to the host galaxy’s mass assumptions. This difference is shown in BI-09 of Table 4, which shows $\Delta_{c/a}$ = 0.01 after 3 Gyr, and $f_{c/a}$ = 1.13. In contrast, the blue curve, representing the $MW_{\text{low}}$ mass model, shows a markedly different behavior. Its axis ratio initially increases before flattening, which could imply a dynamical history distinct from the correct one for the plane of satellites within such a lower mass potential model. Here, for the lower mass Milky Way potential, $\Delta_{c/a}$ = 0.03 after 3 Gyr, and $f_{c/a}$ = 1.27.
@@
The other three panels of Fig. 6 show the effect of both potential models and proper motion uncertainties on the stability of the plane of satellites. In the upper right panel, we have proper motion uncertainties of $\pm$0.04 mas yr-1. The effect under $MW_{\text{fiducial}}$ was already discussed in the previous section. For the $MW_{\text{high}}$ model, the trend is similar to $MW_{\text{fiducial}}$, but with slightly larger axis ratio values. The $MW_{\text{low}}$ model again shows a distinct behavior, where it initially rises and then flattens out over time.
@@
When adding more uncertainties, up to Gaia levels and beyond, the effect strengthens in all the remaining panels, and the plane of the system becomes even wider. Under 0.12 mas yr-1 proper motion uncertainty, all three systems show a similar trend, suggesting that higher proper motion uncertainties overshadow the effect of using an incorrect Milky Way model. This observation indicates that, at higher levels of proper motion uncertainty, the distinctions between the various Milky Way potential models become less pronounced, as the increased uncertainty tends to dominate the inferred dynamics of the satellite plane.
@@
Building upon this, we introduced an additional layer of complexity by incorporating distance uncertainties. A comparison between Fig. 6 and Fig. 7, reveals a noticeable effect of distance uncertainties across all three potential models, but the evolution of the plane does not change drastically. However, in the presence of higher proper motion uncertainties, the effect of distance uncertainties is again overshadowed under all three potential models. Models BI-13 to BI-24 show this behavior when 5% uncertainties are induced.
@@
We also verified that our results remain qualitatively unchanged when using different Milky Way gravitational potentials. Specifically, we performed the analysis with MilkyWayPotential2022 from Gala (Price-Whelan, 2017), including variations in halo mass and proper motion uncertainties, finding no substantial differences in the overall trends (see Appendix B.2).
@@ 3.5 Analysis of individual axes
Figure 8: Comparative analysis of galactic orbital axes under proper motion uncertainties and different potential models at $\theta_{tan}$ = 80°. Panels shows the mean values of 600 major, minor, and intermediate axes calculated through backward integration. Each panel corresponds to a unique combination of proper motion uncertainty and potential models. Dash-dotted curves show the effect of distance uncertainties, while solid curves show results without distance uncertainties. Rows correspond to different levels of proper motion uncertainty, while columns represent the three axes: major (left), intermediate (center), and minor (right).
@@
The behavior of the minor-to-major axial ratio in the $MW_{\text{low}}$ potential model differs from that in other models, warranting a closer examination of how the extent of the test satellite system changes along its axes, from major to minor. Figure 8 illustrates the influence of proper motion uncertainties on the mean values of the three axes: $a$ (major), $b$ (intermediate), and $c$ (minor). The influence of proper motion uncertainties is organized by row, while the axes are displayed by column in the figure. Each panel includes curves in three colors, depicting the impact of different potential models. The solid curves represent the results when $\epsilon_{dist}=0\%$, while the dashed-dotted curves show the results when $\epsilon_{dist}=5\%$. The first row shows $\pm$0.00 mas yr-1 proper motion uncertainties and 0% distance uncertainties. Here, both $MW_{\text{fiducial}}$ and $MW_{\text{high}}$ show similar behavior across the axes, both with and without distance uncertainties. Over time, the axes tend to decrease, particularly under the $MW_{\text{high}}$ potential model. In contrast, $MW_{\text{low}}$ shows a completely different and opposite behavior. During backward integration, all the axes under $MW_{\text{low}}$ increase over time.
@@
In these models, the variations in axis lengths are a direct consequence of changes in gravitational pull due to mass differences. In the $MW_{\text{high}}$ model, increased halo mass leads to more compact axes as the structure becomes more gravitationally bound. Conversely, in the $MW_{\text{low}}$ model, a reduction in mass leads to an elongation of the axes, particularly in $a$ and $b$. This suggests that in the $MW_{\text{low}}$ model, more satellites become unbound, causing them to drift away from their host galaxies. We emphasize that the $MW_{\text{fiducial}}$ model serves as a baseline for the $MW_{\text{high}}$ and $MW_{\text{low}}$ models, with the halo mass adjusted by 40% (increased and decreased, respectively). However, the response of the axis extents is striking, with the Milky Way potential with a 40% reduction in halo mass having a more pronounced effect on the system than a 40% increase. Moreover, including distance errors causes each axis in every potential model to extend slightly, indicating a noticeable, though moderate, effect on the orbital axes.
@@
We also tested and recalculated the $c/a$ ratio by excluding test satellites beyond 300 kpc, considering only those within this threshold, as discussed in Appendix B.3. We find that this does exclusion not result in a major influence.
@@ 3.6 Inferred escaping satellites
Figure 9: Average number of test satellite exceeding a radial distance of 300 kpc under varying proper motion uncertainties. The four panels correspond to different level of proper motion uncertainty. Each panel displays six curves, representing the number of test satellites beyond 300 kpc of distance under three different potential models. Dash-dotted curves show results with distance uncertainties included, while solid curves reflect only proper motion errors.
Figure 9 shows the average number of test satellites that surpass a distance of 300 kpc from the galactic center under proper motion uncertainties and various potential models. The plot is divided into four panels that show three-color curves representing the impact of potential models. The solid curves represent the results for $\epsilon_{dist}=0\%$, while the dashed-dotted curves correspond to $\epsilon_{dist}=5\%$. In the upper left panel, when there are no proper motion uncertainties, the $MW_{\text{low}}$ curve initially shows a relatively stable number of satellites exceeding 300 kpc in backward integration, both with and without distance uncertainties. Over time, the number gradually increases before reaching a plateau. This outcome is expected, as reducing the halo mass by 40% lowers the gravitational binding force, allowing more satellites to become unbound and drift beyond 300 kpc.
Unlike the $MW_{\text{low}}$ model, the other two potential models exhibit opposite behavior. In $MW_{\text{fiducial}}$, the number of test satellites that go beyond 300 kpc decreases over time, and at 5 Gyr, all $N_{\text{sat}}$ test satellites remain under the 300 kpc limit. This trend is even more pronounced in the $MW_{\text{high}}$ potential model, where the number of distant test satellites progressively decreases over time due to the higher halo mass. Furthermore, the other three panels demonstrate the effect of proper motion uncertainties on satellite dynamics, showing that as uncertainties increase, more test satellites remain at distances greater than 300 kpc across all models. Proper motion errors preferentially increase the inferred velocities, and thus satellite energies, effectively ejecting more test satellites.
Furthermore, when comparing the curves with and without distance uncertainties, we observe a clear effect: these uncertainties increase the number of test satellites exceeding 300 kpc. In Table 4, the column $N_{\text{sat}}>$ 300 kpc lists models BI-05 to BI-08 without induced distance errors in $MW_{\text{low}}$, and models BI-17 to BI-20, with distance errors. These results indicate that the inclusion of distance errors at 3 Gyr leads to a greater inferred loss of satellites in the system.